OpenStudy (anonymous):

Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately, how long does it take until the soccer ball hits the ground again?

OpenStudy (dumbcow):

h(t) = -16t^2 +42t = 0

OpenStudy (anonymous):

Um, shouldn't there be a g in there somewhere? 0 = 42-gt, solve for t and double it....?

OpenStudy (anonymous):

yea Im still lost....

OpenStudy (dumbcow):

yes, estudier that is another way to solve it where g=32 ft/sec^2....however artistic use the same equation as before......h(t) = -16t^2 +vt + s, where v = 42 and s = 0.........the ground is when height is zero, so set h(t) = 0 and solve for t

OpenStudy (anonymous):

ok so I got 1.3.... Can someone do me the favor to check and see if Im correct.

OpenStudy (anonymous):

Ah, OK I am used to 9.8, didn't recognize the fps...

OpenStudy (dumbcow):

Artistic, sorry for delay in the reply.......you should get t =2.625........-16t^2 +42t = 0, factor out a t .......-16t+42 = 0 --> 16t = 42 --> t = 42/16 = 2.625